Suppose the mean amount of cholesterol in eggs labeled “large” is 186 milligrams, with standard deviation 7 milligrams. Find the probability that the mean amount of cholesterol in a sample of 144 eggs will be within 1.5 milligrams of the population mean.


1
Expert's answer
2022-04-14T09:41:03-0400

We have a normal distribution, "\u03bc=186,\u03c3=7,n=144,"

"\bar{x}_1=186-1.5=184.5,\\
\bar{x}_2=186+1.5=187.5."

Let's convert it to the standard normal distribution,

"\bar{z}=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}},\\
\bar{z}_1=\cfrac{184.5-186}{7/\sqrt{144}}=-2.57,\\
\bar{z}_2=\cfrac{187.5-186}{7/\sqrt{144}}=2.57,\\
P(184.5<\bar{X}<187.5)=P(-2.57<\bar{Z}<2.57)=\\
=P(\bar{Z}<2.57)-P(\bar{Z}<-2.57)=\\
=0.9949-0.0051=0.9898\text{ (from z-table).}"